OpenStudy (anonymous):

Chemistry question: What is the mass, in grams, of a sample of 6.12 × 1024 atoms of manganese (Mn)? 6.12x10^24 atoms x 1 mol/6.02x10^23 atoms = 10.166 rounded to 10.2 10.2 mol x 54.9380g/1 mol = 558.505 = 559g I was just wondering if I rounded to high to get the right amount of sig figs? Is my answer correct or should I change it?

OpenStudy (anonymous):

I did it again, but I didn't round and this is what I got: 6.12x10^24 atoms x 1 mol/6.02x10^23 atoms = 10.166 10.166 mol x 54.9380g/1 mol = 558.499g

OpenStudy (anonymous):

i think your right i just did the problem and came up with the same answer ;D

OpenStudy (anonymous):

the first one or the second one?

OpenStudy (anonymous):

the second one ;D

OpenStudy (anonymous):

okay, then I'll go with that one! thank you!

OpenStudy (anonymous):

no problemo ;D

OpenStudy (anonymous):

I will say though, I think I might have written the second one incorrectly. For some reason, I din't feel like it's written correctly...

OpenStudy (anonymous):

don't*

OpenStudy (anonymous):

it should be written like this i mole= 588.499g so u wrote it right ;D

OpenStudy (anonymous):

doesn't it have to have 3 sig figs though?

OpenStudy (anonymous):

so would I just change it to 558g?

OpenStudy (anonymous):

im just not sure, I just know my teacher is very strict about sig figs

OpenStudy (anonymous):

yes if u need 3 sig figs then 558 would be correct ;D i was just trying to be accurate sorry!!

OpenStudy (anonymous):

no no, don't be sorry, I like accuracy! lol :) I just wanted to make sure

OpenStudy (anonymous):

Thanks again for ur help!

OpenStudy (anonymous):

uh-huh no problem ;D

OpenStudy (anonymous):

Okay, yeah so just to clarify, my answer is going to be this: 6.12x10^24 atoms x 1 mol/6.02x10^23 atoms = 10.166 10.166 mol x 54.9380g/1 mol = 558.499 = 558g :)

OpenStudy (anonymous):

yea thats right ;D

OpenStudy (anonymous):

okie dokie! :D